OpenStudy (anonymous):

Which point is a solution to the following system of inequalities? x – 2y > 10 3x + 4y < 10 (1, 10) (2, –3) (4, –4) (–2, –5)

OpenStudy (anonymous):

I have to use substitution right?

OpenStudy (anonymous):

Yes, I believe you have to, and if that is the case, the correct point would be (4,-4).

OpenStudy (anonymous):

Can you explain to me how to do it?

OpenStudy (anonymous):

If you substitute the coordinates into their respective variables (x and y) you will see that they fit the inequalities. E.g. (4,-4) for x - 2y > 10. = (4) -2(-4) > 10. = 4 + 8 > 10. = 12>10 (Which is true)

OpenStudy (anonymous):

Is there a specific method which you were taught to use to solve these problems, because it is pretty much trial and error.

OpenStudy (anonymous):

i mean like i can do what you just did but i think there is a way my teacher wanted me to do it but i haven't done it in forever so i dont quite remember

OpenStudy (anonymous):

I would just suggest that you go through the notes you made on this then, or search up the topic on the internet and watch some educational videos for it.

OpenStudy (anonymous):

Otherwise I don't really know how else you would do this, because if there is a method I have probably forgotten it as well.

OpenStudy (anonymous):

Okay thank you(:

OpenStudy (anonymous):

No problem :)